If secA+tanA=p,then what is the value of secA-tanA
Then find the value of these 

Grade:10

2 Answers

Harshit Singh
3 years ago
Dear Student

we know,
sec2A – tan2A=1
using (a-b)(a+b)= a2- b2
(secA-tanA)(secA+tanA)=1
secA+tanA=p
so(secA-tanA)= 1/p
Thanks
Sudhir
3 years ago
Given : secA+tanA= p ...(I)
To find : secA-tanA
 
We know that sec²A – tan²A=1
=> (SecA + TanA )( SecA - TanA) = 1 ..(ii)
Dividing eq.(ii) by eq.(I) We get,
=> Sec A - Tan A = 1/p
 
=>